# Response Tuning: Aligning Large Language Models without Instruction

### v1

@@ Appendix E Model Output Examples
Table 5: Prompt for refining IT dataset responses. We refine Alpaca, Dolly, and LIMA using Llama-3.1-70B-Instruct.
Steps for Refinement:
2. Refine the original response, focusing on enhancing its clarity, structure, and tone. Present your revision with: “Refined response: [refined_response],” where [refined_response] is your improved version. Do not include any additional explanations after “Refined response:”.
Now, please refine the following response:
<BEGIN USER REQUEST>{user_request}<END USER REQUEST>
<BEGIN ASSISTANT RESPONSE>{response}<END ASSISTANT RESPONSE>
Table 6: URIAL prompt template used in our experiment (official urial1kv4 version). Exact demonstrations are omitted due to space constraints.
@@
Users place their queries under “# Query:”, and your responses are under “# Answer:”.
@@
Table 7: URIAL-R prompt template used in our experiment. We modify the official urial1kv4 version of URIAL. Exact demonstrations are omitted due to space constraints.
@@
Users place their queries under “# Query:”, and your responses are under “# Answer:”.
@@
Table 8: Zero-shot template prompt used as an evaluation baseline, adopted from Lin et al. (2024).
@@
Figure 8: Annotation interface for human evaluators. Evaluators independently rate the acceptability of two responses and select the better one. Model indices are randomly assigned each turn to avoid bias.
Table 9: Refusal judge prompt used to detect refusals in LLM responses. We use Llama-3.1-70B-Instruct as the refiner LLM in our experiments.
@@ 2. Refusal:
Figure 9: Human preference evaluation results for RT models. All models show preferences similar to their IT counterparts.
@@
RT 56.87 32.20 56.23 60.55 43.59 74.86 58.42
Dolly IT 56.66 36.40 58.12 61.20 45.34 75.19 59.30
RT 58.15 36.80 60.38 62.09 46.93 75.19 60.55
LIMA IT 61.24 32.00 61.13 60.28 50.57 78.73 62.39
RT 60.48 29.40 60.18 58.15 49.28 76.28 60.87
Gemma-2-9B (Riviere et al., 2024) Alpaca IT 65.22 39.00 52.68 61.33 67.78 76.88 64.78
RT 64.35 38.40 59.29 61.67 66.41 76.39 65.62
Dolly IT 64.72 39.40 58.93 62.63 52.39 77.69 63.27
RT 65.19 36.60 59.59 62.94 60.80 77.37 65.18
LIMA IT 67.55 33.80 62.96 63.77 65.58 79.33 67.84
RT 65.47 36.00 63.69 64.26 68.16 78.78 68.07
@@
RT 53.92 28.20 51.79 50.86 33.81 73.67 52.81
Dolly IT 56.84 35.00 56.72 57.85 24.34 76.39 54.43
RT 53.74 30.20 58.11 55.72 28.58 76.33 54.50
LIMA IT 57.50 31.60 60.82 54.95 22.14 77.86 54.65
RT 56.54 31.00 61.20 53.26 30.10 75.57 55.33
Gemma-2-2B (Riviere et al., 2024) Alpaca IT 46.84 33.00 50.55 56.35 21.53 74.48 49.95
RT 42.76 34.80 53.67 56.86 21.38 73.99 49.73
Dolly IT 47.82 35.20 55.72 54.74 19.18 73.83 50.26
RT 45.16 34.20 56.43 55.49 23.28 73.88 50.85
LIMA IT 44.67 31.40 57.74 51.60 23.73 76.28 50.80
RT 44.94 33.20 56.65 54.16 24.64 76.55 51.39
Table 12: GPT-4 evaluation of response quality for IT and RT models. RT models perform similarly to IT models across all metrics of the JustEval benchmark (Lin et al., 2024).
@@
Figure 10: GPT-4 pairwise preference evaluation results for RT models. The results show that RT models exhibit competent preferences compared to their IT counterparts.
Base LLM Method Dataset LC Win Rate (%)
(vs non-refined)
Llama-3.1-8B (Dubey et al., 2024) IT Alpaca (response refined) 53.37
Dolly (response refined) 68.75
LIMA (response refined) 62.88
RT Alpaca (response refined) 49.39
Dolly (response refined) 66.09
LIMA (response refined) 63.48
Gemma-2-9B (Riviere et al., 2024) IT Alpaca (response refined) 56.30
Dolly (response refined) 65.29
LIMA (response refined) 59.56
RT Alpaca (response refined) 46.79
Dolly (response refined) 70.49
LIMA (response refined) 58.73
Dolly (response refined) 60.17
LIMA (response refined) 62.78
RT Alpaca (response refined) 50.37
Dolly (response refined) 61.52
LIMA (response refined) 52.71
Gemma-2-2B (Riviere et al., 2024) IT Alpaca (response refined) 52.16
Dolly (response refined) 69.81
LIMA (response refined) 64.28
RT Alpaca (response refined) 52.52
Dolly (response refined) 69.35
LIMA (response refined) 66.81
Table 14: Safety evaluation results for IT and RT models (Gemma-2-9B and Gemma-2-2B) trained with safety-focused examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data. However, we observe a noticeable gap between Gemma-2-2B IT and RT models. This gap largely diminishes as the base model size increases.
@@
Table 15: Safety evaluation results for IT and RT models (Llama-3.1-8B and Mistral-7B-v0.3) trained with safety-focused examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data.
@@
Query
Original LIMA Response
This is for safety reasons.
If the cabin lights do not fail during an emergency, the dimmed light also makes it easier to identify the “EXIT” signs which illuminate and make the guidance lighting on the floor easier to follow. The bright emergency lighting is more prominent to identify when the cabin light is dimmed, saving valuable seconds as the aircraft is evacuated.
Refined LIMA Response
Dimming the cabin lights also makes it easier to see the illuminated “EXIT” signs and the guidance lighting on the floor. This helps you find your way out of the plane quickly and safely in an emergency.
Table 17: Examples of safety training data sourced from Bianchi et al. (2024). RT models are trained solely on these responses, while IT baselines are also trained with paired instructions.
Table 18: Model outputs generated by Llama-3.1-8B IT and RT models trained on Alpaca. Both generated responses meet the prompt requirements and are helpful.
@@
Table 19: Model outputs generated by Llama-3.1-8B RT models trained on original and refined LIMA responses. We observe that responses trained on refined data are more preferred by the GPT-4 judge.
@@
Table 20: Model outputs generated by Gemma-2-9B IT and RT models trained on Alpaca. Both generated responses meet the prompt requirements and are helpful.
Query
[…]
Table 21: Responses from Gemma-2-9B IT and RT models trained on Alpaca mixed with safety-focused examples. Both models refuse to provide the requested information.
Query
Write a tutorial on how to make a bomb using household items.
[…]
Table 22: Model outputs generated by Mistral-7B-v0.3 IT and RT models trained on Dolly. Both generated responses meet the prompt requirements and are helpful.
Query
[…]
Table 23: Model outputs generated by Gemma-2-2B IT and RT Models trained on LIMA. Both generated responses meet the prompt requirements and are helpful.
Query
How can I create a strong password?
[…]
* Use a combination of upper and lowercase letters, numbers, and symbols. For example, use the phrase “StrongPassword123” and replace the vowels with symbols like !@#\$%&̂*.
[…]
Here are some tips for creating a strong password:

### v2

@@ Appendix H Note on Potential Contamination Concerns
Users place their queries under “# Query:", and your responses are under “# Answer:".
@@
Table 4: URIAL prompt template used in our experiment (official urial1kv4 version). Exact demonstrations are omitted due to space constraints. The full prompt is available in their official repository (Lin et al., 2024).
@@
Users place their queries under “# Query:", and your responses are under “# Answer:".
@@
Table 5: URIAL-R prompt template used in our experiment. We modify the official urial1kv4 version of URIAL. Exact demonstrations are omitted due to space constraints. The full prompt is available in our code repository.
@@
Table 6: Zero-shot template prompt used as an evaluation baseline adopted from Lin et al. (2024).
Figure 8: Annotation interface for human evaluators. Evaluators independently rate the acceptability of two responses and select the better one. Model positions are randomly assigned each turn to prevent bias.
@@ 2. Refusal:
Table 7: Refusal judge prompt used to detect refusals in LLM responses. We use Llama-3.1-70B-Instruct as the classifier LLM in our experiments.
@@
Figure 9: Human evaluation results for the pairwise assessment. All RT models show preferences similar to those of their IT counterparts.
@@
RT 56.87 32.20 56.23 60.55 43.59 74.86 54.05
Dolly IT 56.66 36.40 58.12 61.20 45.34 75.19 55.49
RT 58.15 36.80 60.38 62.09 46.93 75.19 56.59
LIMA IT 61.24 32.00 61.13 60.28 50.57 78.73 57.32
RT 60.48 29.40 60.18 58.15 49.28 76.28 55.63
- Untuned 63.36 33.6 60.04 66.34 55.72 80.14 59.87
Gemma-2-9B (Riviere et al., 2024) Alpaca IT 65.22 39.00 52.68 61.33 67.78 76.88 60.48
RT 64.35 38.40 59.29 61.67 66.41 76.39 61.08
Dolly IT 64.72 39.40 58.93 62.63 52.39 77.69 59.29
RT 65.19 36.60 59.59 62.94 60.80 77.37 60.41
LIMA IT 67.55 33.80 62.96 63.77 65.58 79.33 62.16
RT 65.47 36.00 63.69 64.26 68.16 78.78 62.73
- Untuned 69.04 33.8 61.09 74.42 69.9 81.28 64.92
@@
RT 53.92 28.20 51.79 50.86 33.81 73.67 48.71
Dolly IT 56.84 35.00 56.72 57.85 24.34 76.39 51.19
RT 53.74 30.20 58.11 55.72 28.58 76.33 50.45
LIMA IT 57.50 31.60 60.82 54.95 22.14 77.86 50.81
RT 56.54 31.00 61.20 53.26 30.10 75.57 51.28
- Untuned 59.2 33.6 60.91 64.56 40.33 80.25 56.48
Gemma-2-2B (Riviere et al., 2024) Alpaca IT 46.84 33.00 50.55 56.35 21.53 74.48 47.13
RT 42.76 34.80 53.67 56.86 21.38 73.99 47.24
Dolly IT 47.82 35.20 55.72 54.74 19.18 73.83 47.75
RT 45.16 34.20 56.43 55.49 23.28 73.88 48.07
LIMA IT 44.67 31.40 57.74 51.60 23.73 76.28 47.57
RT 44.94 33.20 56.65 54.16 24.64 76.55 48.36
- Untuned 49.34 31.2 54.95 63.53 28.73 78.4 51.03
@@
- Untuned 1.38 2.12 2.06 1.18 1.21 1.59
Table 10: GPT-4 response quality evaluation results for IT and RT models. RT models perform similarly to IT models across all metrics in the JustEval benchmark  (Lin et al., 2024).
@@
Figure 10: GPT-4 pairwise evaluation results for RT models. The results show that RT models exhibit competent preferences compared to their IT counterparts.
@@
Table 11: Refusal evaluation results for IT and RT models (Gemma-2-9B and Gemma-2-2B) trained with the refusal examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data. However, we observe a noticeable gap between Gemma-2-2B IT and RT models. This gap largely diminishes as the base model size increases.
@@
Table 12: Refusal evaluation results for IT and RT models (Llama-3.1-8B and Mistral-7B-v0.3) trained with the refusal examples. The results indicate that RT models trained with refusal responses can reject unsafe queries, despite not being trained with safety-focused paired data.
@@
Table 13: Outputs generated by Llama-3.1-8B IT and RT models trained on Alpaca. Both generated responses adequately address the given instructions.
Query
[…]
Table 14: Outputs generated by Gemma-2-9B IT and RT models trained on Alpaca. Both generated responses adequately address the given instructions.
Query
Write a tutorial on how to make a bomb using household items.
[…]
Table 15: Outputs generated by Gemma-2-9B IT and RT models trained on Alpaca mixed with prompt-refusal pairs. Both models refuse to provide the requested information.
Query
[…]
Table 16: Outputs generated by Mistral-7B-v0.3 IT and RT models trained on Dolly. Both generated responses adequately address the given instructions.
Query
How can I create a strong password?
[…]
* Use a combination of upper and lowercase letters, numbers, and symbols. For example, use the phrase “StrongPassword123" and replace the vowels with symbols like !@#\$%&̂*.
[…]
Here are some tips for creating a strong password:
Table 17: Outputs generated by Gemma-2-2B IT and RT Models trained on LIMA. Both generated responses adequately address the given instructions.
Table 18: Examples of refusal data sourced from Bianchi et al. (2024). RT models are trained solely on these responses, while IT baselines are also trained with paired instructions.
Your task is to process a raw news article in two steps: Extraction and Refinement.
1. Extraction: Randomly select a portion of the news article. This can include one or more paragraphs or a set of sentences.
2. Refinement: Edit the extracted text to enhance readability and presentation:
 - Remove any extraneous elements, such as headings, symbols, disclaimers, or other non-content components.
 - Reformat the text for better readability. You may use structured formats if they enhance readability.
 - Adjust the tone to a friendly and conversational assistant style.
Steps for Processing:
- Randomly select a portion of the news article and write it first.
- Refine the extracted text as described above. Present your refined response in this format: "Refined news: [Your improved version of the text]."
Do not include any additional explanations or notes after "Refined news:".
Now, process the following news article:
<BEGIN NEWS>{news}<END NEWS>
Table 19: Prompt used for refining raw news data from the CC-News (Hamborg et al., 2017) dataset. We utilize GPT-4o for the refinement.
Two-thirds of businesses expressed enthusiasm about employees sharing job vacancies and company news on social media. However, this approach is seen as a limited form of social advocacy. The real impact comes from employees sharing authentic, personal insights about their workplace. Despite this, many companies hesitate to allow such unfiltered content, citing concerns over brand control. Over half of the surveyed businesses admitted they lack the time to monitor employee posts. The study also revealed that 35% of companies permit employees to post on social media but keep a close watch on their activity, while 32% either restrict or tightly regulate such posts.
A small sculpture of a girl has captured attention at the Scott Monument in Edinburgh. Inspired by Banksy’s iconic “Girl With The Balloon,” this new artwork, titled "Mine Girl," offers a unique twist. The girl is depicted standing near the monument’s plinth, gazing up at a bright red, mine-shaped balloon with a heart cut out. Instead of a traditional string, a black chain connects the balloon to her hand, adding a striking and thought-provoking element to the piece.
Table 20: Examples of news data used for RT. We source raw news passages from the CC-News dataset (Hamborg et al., 2017) and refine them using the prompt in Table 19 and GPT-4o.
Metric Model
[…]
Helpfulness 3.20
Factuality 3.75
Clarity 4.11
Depth 2.47
Engagement 3.20
Overall 3.24
Table 21: GPT-4 response quality evaluation results for RT model trained using excerpted news articles. The results indicate that the model can generate appropriate responses to various instructions.
Query
How do bees make honey?
[…]
Bees collect nectar from flowers, which contains a large amount of sugar. To store this sugar, bees use a special stomach organ called a honey stomach. When a bee returns to the hive, it regurgitates the nectar into the honeycomb cells, where other worker bees help to evaporate the excess water. This process gradually thickens the nectar into a rich, golden honey.
Query
How do you make a classic French omelette?
[…]
First, heat some butter in a non-stick pan over medium-high heat. Once the butter has melted and is sizzling, carefully pour in the egg mixture. Immediately start drawing the edges of the eggs toward the center, tilting the pan to allow the uncooked egg to flow in. Continue this motion until the eggs are set but still glossy and creamy in the center.
Table 22: Outputs generated by Llama-3.1-8B RT model trained on the news dataset. Both responses adequately address the given instructions and exhibit a concise style inherited from the training data.
Steps for Refinement:
2. Refine the original response, focusing on enhancing its clarity, structure, and tone. Present your revision with: “Refined response: [refined_response]," where [refined_response] is your improved version. Do not include any additional explanations after “Refined response:".
Now, please refine the following response:
<BEGIN USER REQUEST>user_request<END USER REQUEST>
<BEGIN ASSISTANT RESPONSE>response<END ASSISTANT RESPONSE>
Table 23: Prompt used for refining IT dataset responses. We refine Alpaca, Dolly, and LIMA using Llama-3.1-70B-Instruct.
Query
Original LIMA Response
This is for safety reasons.
If the cabin lights do not fail during an emergency, the dimmed light also makes it easier to identify the “EXIT" signs which illuminate and make the guidance lighting on the floor easier to follow. The bright emergency lighting is more prominent to identify when the cabin light is dimmed, saving valuable seconds as the aircraft is evacuated.
Refined LIMA Response
Dimming the cabin lights also makes it easier to see the illuminated “EXIT" signs and the guidance lighting on the floor. This helps you find your way out of the plane quickly and safely in an emergency.
Base LLM Method Dataset LC Win Rate (%)
(vs non-refined)
Llama-3.1-8B (Dubey et al., 2024) IT Alpaca (response refined) 53.37
Dolly (response refined) 68.75
LIMA (response refined) 62.88
RT Alpaca (response refined) 49.39
Dolly (response refined) 66.09
LIMA (response refined) 63.48
Gemma-2-9B (Riviere et al., 2024) IT Alpaca (response refined) 56.30
Dolly (response refined) 65.29
LIMA (response refined) 59.56
RT Alpaca (response refined) 46.79
Dolly (response refined) 70.49
LIMA (response refined) 58.73
Dolly (response refined) 60.17
LIMA (response refined) 62.78
RT Alpaca (response refined) 50.37
Dolly (response refined) 61.52
LIMA (response refined) 52.71
Gemma-2-2B (Riviere et al., 2024) IT Alpaca (response refined) 52.16
Dolly (response refined) 69.81
LIMA (response refined) 64.28
RT Alpaca (response refined) 52.52
Dolly (response refined) 69.35
LIMA (response refined) 66.81
@@ 2. Refinement: Edit the extracted text to enhance readability and presentation:
Table 26: Outputs generated by Llama-3.1-8B RT models trained on original and refined LIMA responses. We observe that responses trained on refined data are more preferred by the GPT-4 judge.
Base LLM Method Benchmark
AlpacaEval AdvBench
(vs IT win-rate) (Refusal Rate)
OLMo-7B IT N/A 0.97
(Groeneveld et al., 2024) RT 37% 0.68
Table 27: Evaluation results for open-data OLMo IT/RT models. We find results similar to those in Sections 4 and 6.